OpenStudy (anonymous):

Given the function f(x) = -2x + 10, find x if f(x) = 18. A 46 B -4 C -26 D -14

OpenStudy (hoblos):

-2x + 10 =18 -2x = 8 x=-4

OpenStudy (anonymous):

Thanks again, hoblos. You're great!

OpenStudy (hoblos):

my pleasure =)